# Classical harmonic analysis viewed through the prism of noncommutative geometry

### v4

@@ 8 Examples of triples and Banach Fredholm modules
By Banach-Alaoglu theorem, there exist $t_{\alpha}\to 0$ and $g_{j}\in\mathrm{L}^{\infty}(\mathbb{C})$ such that $\Delta_{t_{\alpha}}^{(j)}f\overset{}{\longrightarrow}g_{j}$ in the weak* topology of the dual space $\mathrm{L}^{\infty}(\mathbb{C})$. For any $\varphi\in\mathrm{C}_{c}^{\infty}(\mathbb{C})$, we have
$\int_{\mathbb{C}}(\Delta_{t}^{(j)}f)\varphi=\int_{\mathbb{C}}f\frac{\varphi(\cdot-te_{j})-\varphi}{t}\xrightarrow[t\to 0]{}-\int_{\mathbb{C}}f(\partial_{x_{j}}\varphi),$
since $\frac{\varphi(\cdot-te_{j})-\varphi}{t}\to-\partial_{x_{j}}\varphi$ in $\mathrm{L}^{1}(\mathbb{R})$ as $t\to 0$. Passing to the weak* limit gives $\int g_{j}\varphi=-\int f\,\partial_{x_{j}}\varphi$. Thus $g_{j}=\partial_{x_{j}}f$ in $\mathcal{D}^{\prime}$. Consequently $\partial_{x_{j}}f\in\mathrm{L}^{\infty}(\mathbb{C})$. We conclude that $f\in\mathrm{W}^{1,\infty}(\mathbb{C})$.
@@
Proposition 8.27
Suppose that $1<p<\infty$. For any $f\in\mathrm{C}_{c}^{\infty}(\mathbb{C})$ the operator $M_{f}(\mathrm{i}+D)^{-1}$ is compact on the Banach space $\mathrm{L}^{p}(\mathbb{C})\oplus\mathrm{L}^{p}(\mathbb{C})$.
Proof : The resolvent operator $(\mathrm{i}+D)^{-1}\colon\mathrm{L}^{p}(\mathbb{C})\oplus\mathrm{L}^{p}(\mathbb{C})\to\mathrm{W}^{1,p}(\mathbb{C})\oplus\mathrm{W}^{1,p}(\mathbb{C})$ is bounded by definition. Let $f\in\mathrm{C}_{c}^{\infty}(\mathbb{R})$. Consider a bounded open subset $\Omega\subset\mathbb{C}$ with $\mathrm{C}^{1}$ boundary such that $\operatorname{supp}f\subset\Omega$. Note that we have a bounded operator $M_{f}\colon\mathrm{W}^{1,p}(\mathbb{C})\to\mathrm{W}^{1,p}(\Omega)$. We denote by $J\colon\mathrm{W}^{1,p}(\Omega)\to\mathrm{L}^{p}(\Omega)$ the canonical injection, which is compact by [Eva10, Remark p. 289]. If $\mathcal{E}\colon\mathrm{L}^{p}(\Omega)\to\mathrm{L}^{p}(\mathbb{C})$ is the extension map by 0, then the operator $M_{f}(\mathrm{i}+D)^{-1}$ admits the factorization
$M_{f}(\mathrm{i}+D)^{-1}=E\circ J\circ M_{f}\circ(\mathrm{i}+D)^{-1},$
Proposition 8.28
Suppose that $1<p<\infty$. For any function $f\in\mathrm{C}_{c}^{\infty}(\mathbb{C})$ the operator $M_{f}(D+\mathrm{i})^{-1}$ belongs to $S_{\mathrm{app}}^{2,\infty}\big(\mathrm{L}^{p}(\mathbb{C})\oplus\mathrm{L}^{p}(\mathbb{C})\big)$.
Proof : Since $D^{2}=-\Delta\otimes\mathrm{I}_{2}$, we have
$(\mathrm{i}+D)^{-1}=(\mathrm{i}-D)(1+D^{2})^{-1}=(\mathrm{i}-D)(1-\Delta)^{-1}\otimes I_{2}.$
The Bessel operator $(1-\Delta)^{-1}\colon\mathrm{L}^{p}(\mathbb{C})\to\mathrm{W}^{2,p}(\mathbb{C})$ is bounded, while the first order operator $\mathrm{i}-D$ maps $W^{2,p}(\mathbb{C})\oplus\mathrm{W}^{2,p}(\mathbb{C})$ boundedly into $\mathrm{W}^{1,p}(\mathbb{C})\oplus\mathrm{W}^{1,p}(\mathbb{C})$. Set
$S\overset{\mathrm{def}}{=}(\mathrm{i}-D)(1-\Delta)^{-1}\otimes\mathrm{I}_{2}\colon\mathrm{L}^{p}(\mathbb{C})\oplus\mathrm{L}^{p}(\mathbb{C})\to\mathrm{W}^{1,p}(\mathbb{C})\oplus\mathrm{W}^{1,p}(\mathbb{C}).$
Let $f\in\mathrm{C}_{c}^{\infty}(\mathbb{C})$. Fix a closed square $Q\subset\mathbb{C}$ such that $\operatorname{supp}f\subset Q$. Consider the restriction operator
$R_{Q}\colon\mathrm{W}^{1,p}(\mathbb{C})\oplus\mathrm{W}^{1,p}(\mathbb{C})\to\mathrm{W}^{1,p}(Q)\oplus\mathrm{W}^{1,p}(Q)$
and the extension by zero operator $E_{Q}\colon\mathrm{L}^{p}(Q)\oplus\mathrm{L}^{p}(Q)\to\mathrm{L}^{p}(\mathbb{C})\oplus\mathrm{L}^{p}(\mathbb{C})$. Let $J\colon\mathrm{W}^{1,p}(Q)\oplus\mathrm{W}^{1,p}(Q)\hookrightarrow\mathrm{L}^{p}(Q)\oplus\mathrm{L}^{p}(Q)$ be the canonical embedding. Then we have the factorization
(8.48) $M_{f}(\mathrm{i}+D)^{-1}=E_{Q}\underbrace{M_{f}|_{\mathrm{L}^{p}(Q)\oplus\mathrm{L}^{p}(Q)}}_{M_{f}^{Q}}J\underbrace{R_{Q}S}_{S_{Q}}.$
Partition $Q$ into $m^{2}$ pairwise disjoint squares $Q_{1},\dots,Q_{m^{2}}$ of equal area $\frac{|Q|}{m^{2}}$ and side length $\ell_{m}\overset{\mathrm{def}}{=}\frac{\sqrt{|Q|}}{m}$. Consider the average operator $P_{m}\colon\mathrm{W}^{1,p}(Q)\to\mathrm{L}^{p}(Q)$ defined by
$(P_{m}g)(x)\overset{\mathrm{def}}{=}\sum_{k=1}^{m^{2}}\Big(\frac{1}{|Q_{k}|}\int_{Q_{k}}g\Big)1_{Q_{k}}(x),$
and extend componentwise to $\mathcal{P}_{m}\overset{\mathrm{def}}{=}P_{m}\oplus P_{m}\colon\mathrm{W}^{1,p}(Q)\oplus\mathrm{W}^{1,p}(Q)\to\mathrm{L}^{p}(Q)\oplus\mathrm{L}^{p}(Q)$. Clearly, $\operatorname{rank}\mathcal{P}_{m}\leqslant 2m^{2}$. On each square $Q_{k}$, the Poincaré inequality (8.27) yields
[…]
Summing over $k=1,\dots,n$ (the $Q_{k}$ are disjoint) gives
$\left\|g-P_{m}(g)\right\|_{\mathrm{L}^{p}(Q)}^{p}=\sum_{k=1}^{m^{2}}\left\|g-\frac{1}{|Q_{k}|}\int_{I_{k}}g\right\|_{\mathrm{L}^{p}(Q_{k})}^{p}\lesssim\ell_{m}^{p}\sum_{k=1}^{m^{2}}\left\|\nabla g\right\|_{\mathrm{L}^{p}(Q_{k})}^{p}\lesssim\ell_{m}^{p}\left\|\nabla g\right\|_{\mathrm{L}^{p}(Q)}^{p}.$
Taking the $p$-th root, we obtain
$\left\|g-P_{m}(g)\right\|_{\mathrm{L}^{p}(Q)}\leqslant\ell_{m}\left\|\nabla g\right\|_{\mathrm{L}^{p}(Q)}\leqslant\ell_{m}\left\|g\right\|_{\mathrm{W}^{1,p}(Q)}.$
Applying this to each component, we obtain for all $(u,v)\in\mathrm{W}^{1,p}(Q)\oplus\mathrm{W}^{1,p}(Q)$,
$\left\|(J-\mathcal{P}_{m})(u,v)\right\|_{\mathrm{L}^{p}(Q)\oplus\mathrm{L}^{p}(Q)}\leqslant\left\|u-P_{m}u\right\|_{\mathrm{L}^{p}(Q)}+\left\|v-P_{m}v\right\|_{\mathrm{L}^{p}(Q)}$
$\lesssim\ell_{m}\big(\left\|\nabla u\right\|_{\mathrm{L}^{p}(Q)}+\left\|\nabla v\right\|_{\mathrm{L}^{p}(Q)}\big)\lesssim\ell_{m}\left\|(u,v)\right\|_{\mathrm{W}^{1,p}(Q)\oplus\mathrm{W}^{1,p}(Q)}.$
Therefore
(8.49) $\left\|J-\mathcal{P}_{m}\right\|_{\mathrm{W}^{1,p}(Q)\oplus\mathrm{W}^{1,p}(Q)\to\mathrm{L}^{p}(Q)\oplus\mathrm{L}^{p}(Q)}\lesssim\ell_{m}=|Q|^{\frac{1}{2}}m^{-1}.$
Since $\operatorname{rank}\mathcal{P}_{m}\leqslant 2m^{2}$, the definition of approximation numbers gives
[…]
Using the ideal property together with the factorization (8.48), we infer that
[…]
Since $(a_{n})$ is decreasing, for arbitrary $N\geqslant 1$ choose $m=\lfloor\sqrt{(N-1)/2}\rfloor$. Hence $2m^{2}+1\leqslant N$ and we obtain
$a_{N}\big(M_{f}(\mathrm{i}+D)^{-1}\big)\lesssim N^{-\frac{1}{2}}.$
We conclude with (5.9).
The preceding results of this section allow us to state the following corollary.
Corollary 8.29
[…]
Note that we have
[…]
Observe that the function $p\colon\mathbb{C}\to\mathrm{M}_{2}(\mathbb{C})$, $z\mapsto\displaystyle\frac{1}{|z|^{2}+1}\begin{bmatrix}|z|^{2}&z\\ \overline{z}&1\end{bmatrix}$ satisfies $p^{2}=p$ and $\lim_{z\to\infty}p(z)=\begin{bmatrix}1&0\\ 0&0\end{bmatrix}$. That entails that $p-\begin{bmatrix}1&0\\ 0&0\end{bmatrix}$ belongs to the algebra $\mathrm{C}_{0}(\mathbb{C},\mathrm{M}_{2})=\mathrm{M}_{2}(\mathrm{C}_{0}(\mathbb{C}))$. It is well known that the class
$\beta_{\mathbb{C}}\overset{\mathrm{def}}{=}[p]-[1]$
of $\mathrm{K}_{0}(\mathrm{C}_{0}(\mathbb{C}))$ generates the infinite cyclic group $\mathrm{K}_{0}(\mathrm{C}_{0}(\mathbb{C}))\approx\mathbb{Z}$, as discussed in [Eme24, p. 303], [WeO93, p. 104, p. 145, p. 156, p. 177] and [GVF01, p. 122]. This class is called the Bott generator for $\mathbb{C}$. The pairing is described in [CGRS14, pp. 86-86] in the case $p=2$. A similar property holds for the case $p\not=2$.
It is worth noting that by [AIM09, p. 102], the (planar) Beurling-Ahlfors operator $\mathcal{S}\colon\mathrm{L}^{p}(\mathbb{C})\to\mathrm{L}^{p}(\mathbb{C})$ defined by
(8.51) $(\mathcal{S}f)(z)\overset{\mathrm{def}}{=}-\frac{1}{\pi}\operatorname{p.v.}\int_{\mathbb{C}}\frac{f(w)}{(w-z)^{2}}\mathop{}\mathopen{}\mathrm{d}A(w).$
is equal by [AIM09, p. 102] to the square $R^{2}$ of the complex Riesz transform $R$. Moreover, according to [AIM09, Corollary 4.1.1 p. 102], it is a Fourier multiplier with symbol $\big(\frac{\overline{\zeta}}{\zeta}\big)^{2}=\frac{\overline{\zeta}}{|\zeta|}$. This operator has the fundamental property that it turns $\bar{z}$-derivatives into $z$-derivatives, i.e., we have $\mathcal{S}(\partial_{\overline{z}}f)=\partial_{z}f$ for any function $f\in\mathrm{C}^{\infty}_{0}(\mathbb{C})$. Then in this context, we have the following result.
Proposition 8.30
[…]
Proof : Indeed, we have
$\begin{bmatrix}0&\mathcal{S}^{-1}\\ \mathcal{S}&0\\ \end{bmatrix}^{2}=\begin{bmatrix}\mathcal{S}^{-1}\mathcal{S}&0\\ 0&\mathcal{S}\mathcal{S}^{-1}\\ \end{bmatrix}^{2}=\begin{bmatrix}\mathrm{Id}_{\mathrm{L}^{p}(\mathbb{C})}&0\\ 0&\mathrm{Id}_{\mathrm{L}^{p}(\mathbb{C})}\\ \end{bmatrix}$.
Moreover, for any $f\in\mathrm{L}^{\infty}(\mathbb{C}$ we have
According to [AIM09, Theorem 4.6.14 p. 145], for any function $f$ in $\mathrm{VMO}(\mathbb{C})$, the commutator $\mathcal{S},M_{f}]$ is compact on $\mathrm{L}^{p}(\mathbb{C})$. Hence the commutator $\mathcal{S}^{-1},M_{f}]=-\mathcal{S}^{-1}[\mathcal{S},M_{f}]\mathcal{S}^{-1}$ is also compact, as the product of bounded and compact operators. It follows that $\bigg[\begin{bmatrix}0&\mathcal{S}^{-1}\\ \mathcal{S}&0\\ \end{bmatrix},f\bigg]$ is also compact for any $f\in\mathrm{VMO}(\mathbb{C})$.
Remark 8.31
@@
The Dirac operator $D$ is defined in terms of $\gamma$ matrices in direct analogy to commutative tori. Define $N\overset{\mathrm{def}}{=}2^{\lfloor\frac{d}{2}\rfloor}$ and select $N\times N$ complex selfadjoint matrices $\{\gamma_{1},\ldots,\gamma_{d}\}$ satisfying $\gamma_{j}\gamma_{k}+\gamma_{k}\gamma_{j}=2\delta_{j,k}1$. Following [EcI18, (B.6) p. 147] and [GVF01, Definition 12.14 p. 545], we define the unbounded densely defined linear operator
@@
with domain $\ell^{p}_{N}(\mathrm{W}^{1,p}(\mathbb{T}_{\theta}^{d}))$ acting on the complex Hilbert space $\ell^{p}_{N}(\mathrm{L}^{p}(\mathbb{T}^{d}_{\theta}))$. The operator $D$ is selfadjoint. By [GVF01, p. 545], we have
@@ 8.7 Banach Fredholm modules arising from pseudo-differential operators
$\left\|s\right\|_{\mathrm{L}^{p}(M,E)}\overset{\mathrm{def}}{=}\begin{cases}\int_{M}|s(x)|_{E_{x}}^{p}\mathop{}\mathopen{}\mathrm{d}\mu(x)&\text{ if }1\leqslant p<\infty\\ \operatorname*{esssup}_{x\in M}|s(x)|_{E_{x}}&\text{ if }p=\infty\end{cases}.$
@@
Consider an elliptic Hermitian pseudo-differential operator $P\colon\mathrm{C}^{\infty}(M,E)\to\mathrm{C}^{\infty}(M,E)$ of order $m>0$ on a closed Riemannian manifold $M$. Suppose that $1<p<\infty$. Assume that $P$ induces a bisectorial operator on the Banach space $\mathrm{L}^{p}(M,E)$ admitting a bounded $\mathrm{H}^{\infty}(\Sigma^{\mathrm{bi}}_{\theta})$ functional calculus for some angle $0<\theta<\frac{\pi}{2}$. Then $(\mathrm{L}^{p}(M,E),\operatorname{\mathrm{sign}}P)$ is an odd Banach Fredholm module over the algebra $\mathrm{C}(M)$ on the Banach space $\mathrm{L}^{p}(M,E)$.
Proof : By [BaG82b, Proposition 2.4, p. 143] and its proof, the operator $\operatorname{\mathrm{sign}}D\colon\mathrm{L}^{2}(M,E)\to\mathrm{L}^{2}(M,E)$ induces a pseudo-differential operator $\operatorname{\mathrm{sign}}P\colon\mathrm{C}^{\infty}(M)\to\mathrm{C}^{\infty}(M)$ of order 0. Recall that a multiplication operator $M_{f}\colon\mathrm{C}^{\infty}(M)\to\mathrm{C}^{\infty}(M)$ is a differential operator of order 0, hence a pseudo-differential operator of order 0. It is easy to check6665. We warn the reader that the commutator of two pseudo-differential operators on a vector bundle $E$ of order $m$ and $m^{\prime}$ is not necessarily of order $m+m^{\prime}-1$. using the formula [Tre80, (7.2) p. 75] that the commutator $\operatorname{\mathrm{sign}}P,M_{f}]$ is a pseudo-differential operator of order 1-1. Since it has negative order, it induces by Lemma 8.48 a compact operator on the Hilbert space $\mathrm{L}^{2}(M,E)$. Moreover, it equally induces by Theorem 8.49 a bounded operator on the Banach space $\mathrm{L}^{q}(M,E)$ for any $1<q<\infty$. By interpolation [KZPP76, Theorem 3.11 p. 58] relying on the formula (8.77), we conclude that the commutator $\operatorname{\mathrm{sign}}P,M_{f}]$ induces a compact operator on the Banach space $\mathrm{L}^{p}(M,E)$. By approximation, we see that it is true more generally for any function $f\in\mathrm{C}(M)$.
Recall that by [Dieu88, (23.35.2)] the unbounded selfadjoint operator $P$ has compact resolvent on $\mathrm{L}^{2}(M,E)$. Consequently, for any continuous function $\psi\in\mathrm{C}_{0}(\mathbb{R})$ the operator $\psi(P)\colon\mathrm{L}^{2}(M,E)\to\mathrm{L}^{2}(M)$ is compact by [Eme24, Exercise 9.5.4 p. 396]. Since the spectrum of $P$ on $\mathrm{L}^{2}(M,E)$ is discrete, the operator $(\operatorname{\mathrm{sign}}P)^{2}-\mathrm{Id}_{\mathrm{L}^{2}(M,E)}$ can be identified to such an operator $\psi(D)$. Consequently, it is a compact operator on $\mathrm{L}^{2}(M,E)$. Again by interpolation, we obtain that $(\operatorname{\mathrm{sign}}P)^{2}-\mathrm{Id}_{\mathrm{L}^{p}(M,E)}$ is a compact operator on the space $\mathrm{L}^{p}(M,E)$.

### v5

@@ 8.2 The Dirac operator on $\ell^{p}_{N}(\mathrm{L}^{p}(\mathbb{R}^{d}))$
Note that the operator $\partial_{z}(-\Delta)^{-\frac{1}{2}}$ is the complex Riesz transform $R_{\mathbb{C}}\colon\mathrm{L}^{p}(\mathbb{C})\to\mathrm{L}^{p}(\mathbb{C})$ (or complex Hilbert transform) considered in [AIM09, pp. 102-103], [IwG96] and [CDK23]. It is a Fourier multiplier with symbol $\frac{\overline{\zeta}}{|\zeta|}$. By [AIM09, Theorem 4.2.1 p. 103], this operator admits the following integral representation
(8.64) $(R_{\mathbb{C}}f)(z)=\frac{1}{2\pi\mathrm{i}}\operatorname{p.v.}\int_{\mathbb{C}}\frac{\overline{z}-\overline{w}}{|z-w|^{3}}f(w)\mathop{}\mathopen{}\mathrm{d}w=\lim_{\varepsilon\to 0^{+}}\frac{1}{2\pi\mathrm{i}}\int_{w\in\mathbb{C}\setminus B(z,\varepsilon)}\frac{\overline{z}-\overline{w}}{|z-w|^{3}}f(w)\mathop{}\mathopen{}\mathrm{d}w,\quad z\in\mathbb{C}.$
It is known [Dra21, Exercise 1.16 p. 7] that
(8.65) $R_{\mathbb{C}}=R_{2}+\mathrm{i}R_{1},$
where $R_{1},R_{2}\colon\mathrm{L}^{p}(\mathbb{R}^{2})\to\mathrm{L}^{p}(\mathbb{R}^{2})$ are the classical Riesz transforms on $\mathbb{R}^{2}$ described in [Gra14a, Section 5.1.4], which are Fourier multipliers by $\frac{-\mathrm{i}\xi_{1}}{|\xi|}$ and $\frac{-\mathrm{i}\xi_{2}}{|\xi|}$, according to [Gra14a, Proposition 5.1.14]. It is worth noting that by [AIM09, p. 102], the (planar) Beurling-Ahlfors operator $\mathcal{S}\colon\mathrm{L}^{p}(\mathbb{C})\to\mathrm{L}^{p}(\mathbb{C})$ defined by
(8.66) $(\mathcal{S}f)(z)\overset{\mathrm{def}}{=}-\frac{1}{\pi}\operatorname{p.v.}\int_{\mathbb{C}}\frac{f(w)}{(w-z)^{2}}\mathop{}\mathopen{}\mathrm{d}w=-\lim_{\varepsilon\to 0^{+}}\frac{1}{\pi}\int\limits_{w\in\mathbb{C}\setminus B(z,\varepsilon)}\frac{f(w)}{(w-z)^{2}}\mathop{}\mathopen{}\mathrm{d}w,\quad z\in\mathbb{C},$
is equal by [AIM09, p. 102] to the square $R^{2}$ of the complex Riesz transform $R$. Moreover, according to [AIM09, Corollary 4.1.1 p. 102], it is a Fourier multiplier with symbol $\big(\frac{\overline{\zeta}}{\zeta}\big)^{2}=\frac{\overline{\zeta}}{|\zeta|}$ and by [Dra21, p. 7] it is an isometry on the Hilbert space $\mathrm{L}^{2}(\mathbb{C})$ and invertible on the Banach space $\mathrm{L}^{p}(\mathbb{C})$ by [Dra21, Exercise 1.16 p. 8]. By [AIM09, Lemma 4.6.12 p. 142], this operator has the fundamental property that it turns $\bar{z}$-derivatives into $z$-derivatives, i.e., we have
(8.67) $\mathcal{S}(\partial_{\overline{z}}f)=\partial_{z}f.$
for any Schwartz function $f$. We refer to [AIM09] and [Dra21] for more information on this operator. The following result could be deduced from Theorem 8.32 and a locally compact analogue of Proposition 7.7, a statement that we have not yet developed in detail.
Proposition 8.33
[…]
Observe that the function $p_{\mathrm{Bott}}\colon\mathbb{C}\to\mathrm{M}_{2}(\mathbb{C})$, $z\mapsto\displaystyle\frac{1}{|z|^{2}+1}\begin{bmatrix}|z|^{2}&z\\ \overline{z}&1\end{bmatrix}$ satisfies $p_{\mathrm{Bott}}^{2}=p_{\mathrm{Bott}}$ and $\lim_{z\to\infty}p_{\mathrm{Bott}}(z)=\begin{bmatrix}1&0\\ 0&0\end{bmatrix}$. That entails that $p_{\mathrm{Bott}}-\begin{bmatrix}1&0\\ 0&0\end{bmatrix}$ belongs to the algebra $\mathrm{C}_{0}(\mathbb{C},\mathrm{M}_{2})=\mathrm{M}_{2}(\mathrm{C}_{0}(\mathbb{C}))$. It is well known that the class
$\beta_{\mathbb{C}}\overset{\mathrm{def}}{=}[p_{\mathrm{Bott}}]-[1]$
of $\mathrm{K}_{0}(\mathrm{C}_{0}(\mathbb{C}))$ generates the infinite cyclic group $\mathrm{K}_{0}(\mathrm{C}_{0}(\mathbb{C}))\cong\mathbb{Z}$, as discussed in [Eme24, p. 303], [WeO93, p. 104, p. 145, p. 156, p. 177] and [GVF01, p. 122]. This class is called the Bott generator for $\mathbb{C}$. The pairing is described in [CGRS14, pp. 86-86] in the case $p=2$. A similar property holds for the case $p\not=2$.
Then in this context, we also have the following result.
Proposition 8.34
[…]
Proof : Indeed, we have $\begin{bmatrix}0&\mathcal{S}^{-1}\\ \mathcal{S}&0\\ \end{bmatrix}^{2}=\begin{bmatrix}\mathcal{S}^{-1}\mathcal{S}&0\\ 0&\mathcal{S}\mathcal{S}^{-1}\\ \end{bmatrix}^{2}=\begin{bmatrix}\mathrm{Id}_{\mathrm{L}^{p}(\mathbb{C})}&0\\ 0&\mathrm{Id}_{\mathrm{L}^{p}(\mathbb{C})}\\ \end{bmatrix}$. Moreover, for any $f\in\mathrm{C}_{0}(\mathbb{C})$ we have
According to [AIM09, Theorem 4.6.14 p. 145], for any function $f$ in $\mathrm{VMO}(\mathbb{C})$, the commutator $\mathcal{S},M_{f}]$ is compact on $\mathrm{L}^{p}(\mathbb{C})$. Hence the commutator $\mathcal{S}^{-1},M_{f}]=-\mathcal{S}^{-1}[\mathcal{S},M_{f}]\mathcal{S}^{-1}$ is also compact, as the product of bounded and compact operators. It follows that $\bigg[\begin{bmatrix}0&\mathcal{S}^{-1}\\ \mathcal{S}&0\\ \end{bmatrix},f\bigg]$ is also compact for any function $f\in\mathrm{VMO}(\mathbb{C})$. Finally, we have $F\gamma=\begin{bmatrix}0&\mathcal{S}^{-1}\\ \mathcal{S}&0\\ \end{bmatrix}\begin{bmatrix}-\mathrm{Id}&0\\ 0&\mathrm{Id}\\ \end{bmatrix}=\begin{bmatrix}0&\mathcal{S}^{-1}\\ -\mathcal{S}&0\\ \end{bmatrix}=-\begin{bmatrix}-\mathrm{Id}&0\\ 0&\mathrm{Id}\\ \end{bmatrix}\begin{bmatrix}0&\mathcal{S}^{-1}\\ \mathcal{S}&0\\ \end{bmatrix}=-\gamma F$ and the equality $\pi(f),\gamma]=\pi(f)\gamma-\gamma\pi(f)=\begin{bmatrix}M_{f}&0\\ 0&M_{f}\\ \end{bmatrix}\begin{bmatrix}-\mathrm{Id}&0\\ 0&\mathrm{Id}\\ \end{bmatrix}-\begin{bmatrix}-\mathrm{Id}&0\\ 0&\mathrm{Id}\\ \end{bmatrix}\begin{bmatrix}M_{f}&0\\ 0&M_{f}\\ \end{bmatrix}=0$.
Remark 8.35
We will also use the notations $\partial\overset{\mathrm{def}}{=}2\partial_{z}$ and $\partial^{*}\overset{\mathrm{def}}{=}-2\partial_{\overline{z}}$.
@@
The Dirac operator $D$ is defined in terms of $\gamma$ matrices in direct analogy to commutative tori. Define $N\overset{\mathrm{def}}{=}2^{\lfloor\frac{d}{2}\rfloor}$ and select $N\times N$ complex selfadjoint matrices $\gamma_{1},\ldots,\gamma_{d}$ satisfying $\gamma_{j}\gamma_{k}+\gamma_{k}\gamma_{j}=2\delta_{j,k}1$. Following [EcI18, (B.6) p. 147] and [GVF01, Definition 12.14 p. 545], we define the unbounded densely defined linear operator
@@
with domain $\ell^{p}_{N}(\mathrm{W}^{1,p}(\mathbb{T}_{\theta}^{d}))$ acting on the complex Hilbert space $\ell^{p}_{N}(\mathrm{L}^{p}(\mathbb{T}^{d}_{\theta}))$. The operator $D$ is selfadjoint if $p=2$. By [GVF01, p. 545], we have
@@ 8.8 Banach Fredholm modules arising from pseudo-differential operators
(8.96) $\left\|s\right\|_{\mathrm{L}^{p}(M,E)}\overset{\mathrm{def}}{=}\begin{cases}\big(\int_{M}|s(x)|_{E_{x}}^{p}\mathop{}\mathopen{}\mathrm{d}\mu(x)\big)^{\frac{1}{p}}&\text{ if }1\leqslant p<\infty\\ \operatorname*{esssup}_{x\in M}|s(x)|_{E_{x}}&\text{ if }p=\infty\end{cases}.$
@@
Consider an elliptic Hermitian pseudo-differential operator $P\colon\mathrm{C}^{\infty}(M,E)\to\mathrm{C}^{\infty}(M,E)$ of order $m>0$ on a closed Riemannian manifold $M$. Suppose that $1<p<\infty$. Assume that $P$ induces a bisectorial operator on the Banach space $\mathrm{L}^{p}(M,E)$ admitting a bounded $\mathrm{H}^{\infty}(\Sigma^{\mathrm{bi}}_{\theta})$ functional calculus for some angle $0<\theta<\frac{\pi}{2}$. Then $(\mathrm{L}^{p}(M,E),\mathrm{sign}P)$ is an odd Banach Fredholm module over the algebra $\mathrm{C}(M)$ on the Banach space $\mathrm{L}^{p}(M,E)$.
Proof : By [BaG82b, Proposition 2.4, p. 143] and its proof, the operator $\mathrm{sign}D\colon\mathrm{L}^{2}(M,E)\to\mathrm{L}^{2}(M,E)$ induces a pseudo-differential operator $\mathrm{sign}P\colon\mathrm{C}^{\infty}(M)\to\mathrm{C}^{\infty}(M)$ of order 0. Recall that a multiplication operator $M_{f}\colon\mathrm{C}^{\infty}(M)\to\mathrm{C}^{\infty}(M)$ is a differential operator of order 0, hence a pseudo-differential operator of order 0. It is easy to check6665. We warn the reader that the commutator of two pseudo-differential operators on a vector bundle $E$ of order $m$ and $m^{\prime}$ is not necessarily of order $m+m^{\prime}-1$. using the formula [Tre80, (7.2) p. 75] that the commutator $\mathrm{sign}P,M_{f}]$ is a pseudo-differential operator of order 1-1. Since it has negative order, it induces by Lemma 8.54 a compact operator on the Hilbert space $\mathrm{L}^{2}(M,E)$. Moreover, it equally induces by Theorem 8.55 a bounded operator on the Banach space $\mathrm{L}^{q}(M,E)$ for any $1<q<\infty$. By interpolation [KZPP76, Theorem 3.11 p. 58] relying on the formula (8.97), we conclude that the commutator $\mathrm{sign}P,M_{f}]$ induces a compact operator on the Banach space $\mathrm{L}^{p}(M,E)$. By approximation, we see that it is true more generally for any function $f\in\mathrm{C}(M)$.
Recall that by [Dieu88, (23.35.2)] the unbounded selfadjoint operator $P$ has compact resolvent on $\mathrm{L}^{2}(M,E)$. Consequently, for any continuous function $\psi\in\mathrm{C}_{0}(\mathbb{R})$ the operator $\psi(P)\colon\mathrm{L}^{2}(M,E)\to\mathrm{L}^{2}(M)$ is compact by [Eme24, Exercise 9.5.4 p. 396]. Since the spectrum of $P$ on $\mathrm{L}^{2}(M,E)$ is discrete, the operator $(\mathrm{sign}P)^{2}-\mathrm{Id}_{\mathrm{L}^{2}(M,E)}$ can be identified to such an operator $\psi(D)$. Consequently, it is a compact operator on $\mathrm{L}^{2}(M,E)$. Again by interpolation, we obtain that $(\mathrm{sign}P)^{2}-\mathrm{Id}_{\mathrm{L}^{p}(M,E)}$ is a compact operator on the space $\mathrm{L}^{p}(M,E)$.